OpenStudy (anonymous):

Find the solution to the system of equations. x + y + z = 8 y + z = 5 z = 2 A. x = 3, y = 3, z = 2 B. x = 3, y = 3, z = 1 C. x = 8, y = 3, z = 2 D. x = 4, y = 3, z = 2

OpenStudy (anonymous):

It's A

OpenStudy (anonymous):

Use z value to find y Use z and y value to find x It cannot be B, because z must equal 2

OpenStudy (anonymous):

because you sub z = 2 into y + z = 5 to get y +2 = 5 y = 5-2 y = 3 now u have both values of z and y so just sub them into the first one x + y + z = 8 x + 3 + 2 = 8 x + 5 = 8 x = 8-5 x = 3 so therefore answer is A x=3, y=3 and z=2

OpenStudy (anonymous):

thanks you guys